OpenStudy (anonymous):

Solve the following system of equations. Enter the x-coordinate of the solution. Round your answer to the nearest tenth. 5x+2y=22 -2x+6y=3

OpenStudy (anonymous):

3.71 i think

OpenStudy (hockeychick23):

3(5x+2y=22)--> 15x+6y=66 -2x+6y=3----> -2x+6y=3 17x=63 x=3.88235294 or 63/17